# On the interaction of dilatancy and friction in the behavior of fluid–saturated sheared granular materials: a coupled Computational Fluid Dynamics–Discrete Element Method study

### v1

@@ II.2.1 Model setup and particle generation
The simulation begins with the generation of particles. As illustrated in Fig. 1, particles are poured into a simulation box, allowed to settle into equilibrium through cycling, and then trimmed to a specified height to form the final granular column as shown in Fig. 1(a). Each column consists of a polydisperse packing of particles, drawn from a normal Gaussian distribution with a mean diameter $d_{p}$ and standard deviation $0.1d_{p}$. The corresponding grain size distribution is presented in Fig. 1(b). To isolate the effects of initial packing on failure and flow behavior under both dry and fluid-saturated conditions, all columns are prepared with identical geometry: an initial length of $L_{i}=62.5d_{p}$, height $H_{i}=50d_{p}$, and width $L_{z}=25d_{p}$. The influence of aspect ratio on flow dynamics has been addressed in previous work Jing et al. (2018). The computational domain, which houses the column, is a rectangular box with dimensions $L_{x}=261.25d_{p}$, $L_{y}=71d_{p}$, and $L_{z}=25d_{p}$. For fluid-saturated cases, this domain is entirely filled with fluid.
@@
Figure 1: Model setup for subaerial and subaqueous conditions: (a) grains configuration, (b) grain size distribution, (c) first step of the model setup, where grains are poured into the simulation box and allowed to settle (initial height, $H_{i}=50d_{p}$ or 0.05 m, and initial length, $L_{i}=62.5d_{p}$ or 0.06 m), (d) second step of the model setup, where failure is induced by removing the supporting wall, (e) initial state of the grains in subaqueous conditions before failure (length, $L_{x}=261.25d_{p}$ or 0.23 m; height, $L_{y}=71d_{p}$ or 0.071 m; and width, $L_{z}=25d_{p}$ or 0.025 m of the computational domain). Non-slip boundary conditions with frictional walls having the same coefficient of friction (0.5) as the grains, and (f) final state of the grains in subaqueous conditions after failure (final height, $H_{f}$, and final length, $L_{f}$, of the deposit).
Boundary conditions in the CFD and DEM domains are imposed to replicate a quasi-two-dimensional configuration. In CFD, no-slip conditions are applied on all walls, including the base. For DEM, the boundaries are frictional, with the wall–particle friction coefficient matching that of the grains. Although the DEM framework accommodates a variety of particle shapes and sizes, this study focuses on spherical grains. Relevant particle and fluid properties are provided in Table 1. DEM parameters include a grain density $\rho_{p}=2500.0\,\text{kg/m}^{3}$, Poisson’s ratio $\nu=0.23$, and coefficient of restitution $e=0.2$. The CFD mesh consists of cells with dimensions $9d_{p}\times 6d_{p}\times 3d_{p}$. A time step of $\Delta t=10^{-7}\,\text{s}$ ensures tight coupling between the solvers via iterative data exchange. Simulations start by initializing both fluid and granular fields. The DEM solver updates grain positions and velocities using fine-scale time integration, resolving collisions and contact mechanics. Simultaneously, interaction forces, such as drag and pressure, are computed from relative motion and volume fractions. These forces are used to update the fluid momentum field, while the revised fluid properties influence subsequent grain dynamics. This loop is executed at each time step, maintaining consistency in the fluid–particle interaction.
@@
In the second step, particles are packed into the simulation domain. The chosen packing method can significantly influence the resulting packing fraction. In this study, inlet velocity was varied, and grains were allowed to settle until reaching a stable configuration, as shown in Fig. 1(c). The packing fraction ($\phi$), defined in Eq. 10 as the proportion of the total volume occupied by particles, depends on the number of particles ($N_{\text{particle}}$), the volume of an individual particle ($V_{\text{particle}}$), and the volume of the unit cell containing the particles ($V_{\text{unit cell}}$). It can be adjusted by modifying particle size, shape, or the packing method. For example, introducing smaller particles into the voids of a simple cubic packing can increase $\phi$. Using this approach, two packing fractions were achieved in this study: 0.579 (dense model) and 0.444 (loose model).
@@
Once the desired packing fraction is achieved, the DEM simulation allows the grains to settle to rest. After settling, the supporting wall is removed to simulate the perturbation or failure of the landslide, as shown in Fig. 1(d). The simulation then computes the forces and movements of the particles based on their interactions with each other and with their environment. Figs. 1(e) and 1(f) illustrate the state and movement of the granular material before and after the failure, respectively.
@@ II.2.2 Coarse-graining
[…]
By minimizing the sum of squared residuals between the observed and predicted velocities described by Eq. 13, we determine both the average velocity and the velocity gradient tensor $L_{ij}$ in each RVE. The symmetric part of $L_{ij}$ yields the deformation rate tensor $D_{ij}$, as defined in Eq. 14. This gives us the components of the strain rate tensor ${D}$. Similarly, the stress tensor $T_{ij}$ is computed using Eq. 15, which accounts for all interparticle contact forces within a cell. From this, we extract the normal and shear stress components relevant to local material response.
The variables used in these equations are defined as follows: $x_{i}$ and $X_{j}$ are the deformed and reference coordinates, respectively; $x_{0i}$ is a constant translation vector; $v_{i}$ and $v^{0}_{i}$ are the components of the instantaneous and constant velocity vectors; $L_{ij}$ and $L_{ji}$ represent the components of the velocity gradient tensor and its transpose, respectively; $D_{ij}$ is the component of the symmetric deformation rate tensor; $V$ is the RVE cell volume; $N$ is the number of particles in the cell; $N_{l}$ is the number of contacts for particle $l$; $\Delta x_{i}^{(k,l)}$ is the $i$-th component of the vector from the center of particle $l$ to its $k$-th contact; and $F_{j}^{(k,l)}$ is the $j$-th component of the force on particle $l$ from its $k$-th contact.
@@
To explore the nature of failure and flow, dimensionless numbers, specifically the viscous and inertial numbers, are used to distinguish between flow regimes by comparing competing physical forces. The viscous number, $I_{v}$, quantifies the relative importance of viscous forces to confining pressure, as defined in Eq. 16. It is commonly applied in systems where fluid resistance dominates particle dynamics, such as dense suspensions or saturated soils. Larger values of $I_{v}$ correspond to regimes where viscous effects are more significant. In contrast, the inertial number, $I_{n}$, characterizes the ratio of inertial to confining forces in dry or rapid granular flows, where particle collisions and momentum exchange dominate, as given by Eq. 16. Higher values of $I_{n}$ indicate stronger inertial effects. In these expressions, $\eta$ is the dynamic viscosity of the fluid, $\dot{\gamma}$ is the shear strain rate, $P$ is the confining pressure, $d$ is the particle diameter, and $\rho_{s}$ is the particle density.
$I_{v}=\frac{\eta\dot{\gamma}}{P}\quad\text{and}\quad I_{n}=\frac{\dot{\gamma}d}{\sqrt{P/\rho_{s}}}$ (16)
@@ III Mechanism of failure
The failure mechanisms of granular columns differ significantly between subaerial and subaqueous conditions, with the initial packing fraction playing a crucial role in these behaviors. Subaqueous systems exhibit reduced runout distances and slower failure for both dense and loose configurations due to fluid interactions, with dense models being the most stable, a result consistent with previous studies Zhang et al. (2020); Montellà et al. (2023). In contrast, subaerial systems fail more rapidly and exhibit greater sensitivity to packing fraction Nguyen, Bui, and Fukagawa (2015). Loose models undergo more pronounced and extended collapses, while dense models fail abruptly but with limited spreading. Fig. 2 compares velocity profiles at two time steps (0.04 s and 0.4 s) to illustrate the contrasting failure and runout behaviors across dense and loose configurations in both environments. In subaqueous systems, fluid interactions, such as drag and pore pressure feedback, moderate failure dynamics, leading to slower progression and reduced energy dissipation compared to subaerial cases, although final runout distances may be similar in some cases Zhang et al. (2020).
@@
The dense subaqueous model (Fig. 2(a) and Fig. 2(b)) shows minimal runout, with a maximum of 0.054 m, which is approximately 61% shorter than the maximum runout observed in subaerial conditions. The loose subaqueous model (Fig. 2(c) and Fig. 2(d)) exhibits a slightly longer runout of 0.062 m, yet still 59% shorter than its subaerial counterpart. These reductions underscore the dampening influence of fluid on collapse extent. As shearing initiates, dilation occurs due to interlocking grains pushing against one another, increasing the pore pressure and mobilizing peak shear resistance. This interparticle interlocking elevates the internal friction angle and temporarily raises shear strength Montellà et al. (2023). Following this peak, the friction angle decreases, and failure progresses gradually with limited velocity magnitudes throughout the collapse.
Figure 2: Comparison of failure and runout dynamics using velocity for subaqueous and subaerial models with dense ($\phi$ = 0.579) and loose ($\phi$ = 0.444) configurations at two time steps: 0.04 s and 0.4 s. Subaqueous models exhibit slower failure and reduced runout distances compared to subaerial models, due to fluid interactions. Panels (a) and (b) show the dense subaqueous model with a minimal runout of 0.054 m (61% shorter than the maximum), while (c) and (d) represent the loose subaqueous model, which displays greater displacement and a more extended runout of 0.062 m (59% shorter). Panels (e) and (f) display the dense subaerial model, characterized by abrupt failure with a runout of 0.15 m (maximum), and (g) and (h) illustrate the loose subaerial model, which also exhibits a runout of 0.15 m (maximum).
@@
Figure 3: Depiction of shear zone development and pore pressure evolution at two time steps (0.1 s and 0.8 s) for dense ($\phi$=0.579) and loose ($\phi$=0.444) subaqueous models. Panels (a) and (b) show shear zones, defined by zero volumetric strain with the critical state framework Wood (1991), for the dense model at the two respective time steps, while panels (e) and (f) show the corresponding results for the loose model. Panels (c) and (d) illustrate pore pressure changes in the dense model, which begins with lower initial values that increase over time; in contrast, panels (g) and (h) show the loose model, which starts with higher initial pore pressures that decrease with continued deformation
In contrast, the loose subaqueous model experiences greater displacement, as shown in Fig. 2(c) and Fig. 2(d). The particles in this configuration are initially loosely packed, resulting in a lower packing fraction and more intergranular void space. This allows for particle rearrangement and contraction during shearing. As shearing progresses, contraction reduces the overall volume and diminishes shear strength due to the loss of effective interlocking, an effect well documented in subaqueous and analogous slope failure studies Zhang and Puzrin (2021); Carey et al. (2019). This internal reorganization leads to more pronounced motion and a longer runout compared to the dense model. Experimental and numerical studies have further shown that loose granular configurations, which are highly sensitive to pore pressure, exhibit increased flow mobility under submerged conditions Lee and Chen (2022). The loose subaqueous model also displays higher velocity, reflecting a more dynamic collapse process. Overall, the contrast between the dense and loose subaqueous models highlights the failure behavior governed by the critical role of initial packing density and fluid–particle interactions. While the dense model benefits from shear-induced dilation and associated frictional strengthening, the loose model, dominated by contraction, increased pore pressure, and reduced interparticle interlocking, undergoes a more mobile and extensive failure.
In subaerial systems, where fluid effects are absent, failure mechanisms are primarily governed by the initial packing fraction of the granular material. The dense model, illustrated in Fig. 2(e) and Fig. 2(f), demonstrates failure when internal stresses exceed the material’s strength. Although the collapse is dilative, meaning that volume increases as particles are sheared, the runout distance remains comparable to that of the loose subaerial case. This limited difference arises from the high initial packing fraction, which enhances interlocking and resistance to deformation Tapia et al. (2013). In the loose subaerial model, shown in Fig. 2(g) and Fig. 2(h), exhibits runout, extending to 0.15 meters. The lower initial packing fraction allows particles to rearrange more readily and contract during shearing, leading to increased mobility, which is comparable to the dense model. This contraction reduces shear strength and promotes more extensive displacement, as loose grains are less able to resist movement. Experimental findings confirm that such configurations lead to faster and broader collapses, emphasizing the importance of initial packing in governing failure dynamics under dry conditions Bougouin and Lacaze (2018a). Consequently, the contrast between dense and loose subaerial models underscores the critical role of granular structure in determining collapse behavior in the absence of fluid interactions.
There are contrasting behaviors between dense and loose subaqueous models in terms of shear zone evolution and pore pressure dynamics. Dense specimens exhibit a slower, progressive build-up of pore pressure under undrained cyclic loading, attributed to their higher resistance to deformation and reduced compressibility at greater densities Baćić and Herle (2023). In contrast, loose models show broader, more diffuse shear zones accompanied by decreasing pore pressure, as fluid redistributes during contraction and particle rearrangement. This behavior is well captured in multiphase modeling studies, which demonstrate that shear-induced volume changes in loosely packed columns lead to distinct pore-pressure feedback mechanisms, influencing both the timing and extent of collapse Lee (2020). Experimental studies further underscore the role of fluid: under saturated conditions, reduced flow mobility and increased resistance promote denser frontal accumulation Tong, Qu, and He (2022). Fig. 3 provides a detailed visualization of these behaviors at two distinct time steps, 0.1 seconds and 0.8 seconds, highlighting the influence of initial packing density on deformation patterns and pore pressure evolution in fluid-saturated granular columns.
Shear zones in the dense subaqueous model correspond to localized regions of deformation where volumetric strain approaches zero, an interpretation consistent with the critical state framework Wood (1991). At the early time step of 0.1 seconds, Fig. 3(a), these zones initiate near the base and extend upward in a concentrated pattern. By 0.8 seconds, Fig. 3(b), the shear zones become more pronounced and well-defined, reflecting the progressive nature of shearing during failure. The high initial packing fraction in dense granular configurations enhances frictional resistance, limits deformation, and promotes the development of narrower, more stable shear zones over time Baćić and Herle (2023); Desrues and Viggiani (2004). These features align with models of dilative deformation in dense, low-permeability granular media, where pore pressure increases due to particle interlocking and restricted fluid redistribution Petford and Koenders (2003). Initially, pore pressure remains low in the dense subaqueous model because the tightly packed structure inhibits fluid mobility, as shown in Fig. 3(c). As shearing progresses, Fig. 3(d), pore pressure gradually increases, reaching higher values by 0.8 seconds. This rise results from excess pressure generation driven by shearing-induced dilation and the limited ability of fluids to dissipate, characteristics observed in compacted, low-permeability sediments in subduction zones Nikolinakou et al. (2023). The panels illustrate how pore pressure intensifies in tandem with the development and localization of shear zones.
The shear zones in the loose subaqueous model exhibit broader and more diffuse deformation patterns compared to those in the dense model. At 0.1 seconds, Fig. 3(e), these zones appear wider and less sharply defined, reflecting the lower initial packing fraction, which promotes particle rearrangement and volumetric contraction. By 0.8 seconds, Fig. 3(f), the deformation zones expand further, illustrating the reduced shear resistance and enhanced mobility in loosely packed configurations. This behavior is consistent with two-phase flow simulations, which show that low-density granular columns develop widespread shear zones early during collapse, accompanied by elevated pore pressures that dissipate as the material contracts Lee (2020). In such systems, shear bands propagate more readily due to higher initial pore pressures and lower effective stresses. As deformation progresses, these broad zones continue to widen, underscoring the role of initial packing fraction in controlling the kinematic spread of failure.
In the loose subaqueous model, pore pressure initially reaches elevated levels, Fig. 3(g), because the open grain framework allows greater fluid saturation. As shearing advances, Fig. 3(h), pore pressure decreases due to fluid drainage facilitated by contraction and increased permeability behavior that aligns with analytical models describing pressure evolution in contractant granular media Petford and Koenders (2003). This reduction in pore pressure supports ongoing deformation and further expansion of the shear zones by illustrating how pore pressure diminishes over time as particles reorganize and the granular structure evolves under submerged conditions.

### v2

@@ II.2.1 Model setup and particle generation
The simulation begins with the generation of particles. As illustrated in Fig. 1, particles are poured into a simulation box, allowed to settle into equilibrium through cycling, and then trimmed to a specified height to form the final granular column, as shown in Fig. 1(a). Each column consists of a polydisperse packing of particles, drawn from a normal Gaussian distribution with a mean diameter $d_{p}$ and a standard deviation $0.1d_{p}$. The corresponding grain size distribution is presented in Fig. 1(b). To isolate the effects of initial packing on failure and flow behavior under both dry and fluid-saturated conditions, all columns are prepared with identical geometry: an initial length of $L_{i}=62.5d_{p}$, a height $H_{i}=50d_{p}$, and a width $L_{z}=25d_{p}$. The influence of aspect ratio on flow dynamics has been addressed in previous work Jing2018. The computational domain, which houses the column, is a rectangular box with dimensions $L_{x}=261.25d_{p}$, $L_{y}=71d_{p}$, and $L_{z}=25d_{p}$. For fluid-saturated cases, this domain is entirely filled with fluid.
@@
Figure 1: Model setup for subaerial and subaqueous conditions: (a) grain configuration, (b) grain size distribution, (c) first step of the model setup, where grains are poured into the simulation box and allowed to settle (initial height, $H_{i}=50d_{p}$ or 0.05 m, and initial length, $L_{i}=62.5d_{p}$ or 0.06 m), with aspect ratio, A = $H_{i}$$L_{i}$ = 0.8, (d) second step of the model setup, where failure is induced by removing the supporting wall, (e) initial state of the grains in subaqueous conditions before failure (length, $L_{x}=261.25d_{p}$ or 0.23 m; height, $L_{y}=71d_{p}$ or 0.071 m; and width, $L_{z}=25d_{p}$ or 0.025 m of the computational domain). Non-slip boundary conditions with frictional walls having the same coefficient of friction (0.5) as the grains, and (f) final state of the grains in subaqueous conditions after failure (final height, $H_{f}$, and final length, $L_{f}$, of the deposit).
Boundary conditions in the CFD and DEM domains are imposed to replicate a quasi-two-dimensional configuration. In CFD, no-slip conditions are applied on all walls, including the base. For DEM, the boundaries are frictional, with the wall–particle friction coefficient matching that of the grains. Although the DEM framework accommodates a variety of particle shapes and sizes, this study focuses on spherical grains. Relevant particle and fluid properties are provided in Table 1. DEM parameters include a grain density $\rho_{p}=2500.0\,\text{kg/m}^{3}$, Poisson’s ratio $\nu=0.23$, and coefficient of restitution $e=0.2$. Grain–grain and wall–grain Coulomb friction were fixed at $\mu=\mu_{w}=0.5$, a representative value for quartz sands. In this work, our primary target is to investigate poromechanical coupling; a systematic variation of the grain-grain friction coefficient, $\mu$, and its effect on the failure style and rheology may be examined in future investigations.
@@
The CFD mesh consists of cells with dimensions $9d_{p}\times 6d_{p}\times 3d_{p}$. Cells were $9d_{p}\times 6d_{p}\times 3d_{p}$ (aspect ratio 3). The minimum grid spacing of $3d_{p}$ is consistent with unresolved CFD–DEM grid-to-particle sizing guidelines of $\sim 2$$4d_{p}$ established in verification studies WangLiu2021; Andrews1996. A time step of $\Delta t=10^{-7}\,\text{s}$ ensures tight coupling between the solvers via iterative data exchange. A time step of $\Delta t=10^{-7}\,\mathrm{s}$ is the fluid step; MFIX then subcycles the DEM according to the Hertzian criterion, $\Delta t_{\mathrm{DEM}}=t_{\mathrm{coll}}/N_{\mathrm{coll}}$, where $t_{\mathrm{coll}}$ is computed from the restitution–damping–stiffness relations, Eqs. 31–33, 36–37 using the MFIX–DEM framework for multiphase particulate flows MFIXDEM2012 and depends on $(E,\nu,\rho_{p},d_{p},e)$. In our model, $E=10$ GPa, $\nu=0.23$, $\rho_{p}=2500~\mathrm{kg\,m^{-3}}$, $\bar{d}_{p}=1$ mm, and $e=0.2$, giving $\Delta t_{\mathrm{DEM}}\approx 10^{-8}\text{–}10^{-7}$ s. Simulations start by initializing both fluid and granular fields. The DEM solver updates grain positions and velocities using fine-scale time integration, resolving collisions and contact mechanics. Simultaneously, interaction forces, such as drag and pressure, are computed from relative motion and volume fractions. These forces are used to update the fluid momentum field, while the revised fluid properties influence subsequent grain dynamics. This loop is executed at each time step, maintaining consistency in the fluid–particle interaction. With the grid and time–stepping choices specified above, the outputs were insensitive to modest mesh coarsening/refinement around the previously stated $3d_{p}$ minimum and to halving/doubling both the DEM subcycling factor and the fluid step, consistent with MFIX–DEM verification and validation guidance (MFIXDEM2012; MFIX_DEM_Ref) and time–integration recommendations (MFIX_Time).
In the second step, particles are packed into the simulation domain. The chosen packing method can significantly influence the resulting packing fraction. In this study, the inlet velocity was varied, and grains were allowed to settle until reaching a stable configuration, as shown in Fig. 1(c). The packing fraction ($\phi$), defined in Eq. 12 as the proportion of the total volume occupied by particles, depends on the number of particles ($N_{\text{particle}}$), the volume of an individual particle ($V_{\text{particle}}$), and the volume of the unit cell containing the particles ($V_{\text{unit cell}}$). It can be adjusted by modifying particle size, shape, or the packing method. For example, introducing smaller particles into the voids of a simple cubic packing can increase $\phi$. Using this approach, two packing fractions were achieved in this study: 0.579 (dense configuration) and 0.444 (loose configuration). To isolate preparation effects from size-distribution effects, we fixed a narrow, truncated-normal particle-size distribution (PSD) ($\sigma_{d}\approx 0.1\,\bar{d}_{p}$). Thus, the dense ($\phi=0.579$) and loose ($\phi=0.444$) states reported here arise primarily from the inlet-velocity compaction protocol rather than differences in polydispersity. We denote $d_{p}$ as particle diameter, with $\bar{d}_{p}$ and $d_{50}$ its mean and median. Similarly, it also limits the increase in computational time that would occur for a few tens of thousands of grains if a log-normal distribution were used Shire2020, and avoids introducing small grains for which cohesive forces become significant CohesiveGranularMaterial2025. Building on this framework by incorporating wider log-normal distributions would be a valuable direction for future investigations.
@@
Once the desired packing fraction is achieved, the DEM simulation allows the grains to settle to rest. After settling, the supporting wall is removed to simulate the perturbation or the failure of the landslide, as shown in Fig. 1(d). The simulation then computes the forces and movements of the particles based on their interactions with each other and with their environment. Figs. 1(e) and 1(f) illustrate the state and movement of the granular material before and after the failure, respectively.
@@ II.2.2 Coarse-graining
We use coarse-graining techniques to transition from the microscale to the mesoscale, employing a least-squares regression to compute the velocity gradient tensor L{L} within a Representative Volume Element (RVE) cell, based on the positions and velocities of particles Rycroft2009. Granular elements are defined using large-scale, three-dimensional Discrete Element Method (DEM) simulations. The size of the RVE is chosen to match the scale of dynamical correlations in dense granular flows, which emerge over displacements of a few particle diameters as particles transition from superdiffusive to diffusive behavior Choi2004, and are typically on the order of 3 to 5 particle diameters Rycroft2006. By analyzing the simulation as an ensemble of deforming granular elements, we can locally track material evolution. This framework provides an approximate connection between continuum theories and particle-based simulations, facilitating the study of local relationships between normalized deformation rate, packing fraction, apparent friction, pore pressure, and strain. The input data consist of instantaneous particle positions x{x}, given by Eq. 13, and velocities v{v}, given by Eq. 15, for all particles within the RVE. The domain is discretized into RVE cells in $(x,y,z)$, and average quantities are computed per bin. For each bin, the deformation gradient tensor $\mathbf{F}_{ij}$ is calculated as defined in Eq. 14. From $\mathbf{F}_{ij}$, additional kinematic measures such as the right stretch tensor and strain tensors, Cauchy-Green, Green, and Hencky, are computed.
By minimizing the sum of squared residuals between the observed and predicted velocities as described by Eq. 15, we determine both the average velocity and the velocity gradient tensor $L_{ij}$ in each RVE. The symmetric part of $L_{ij}$ yields the deformation rate tensor $D_{ij}$, as defined in Eq. 16. This gives us the components of the strain rate tensor D{D}. Similarly, the stress tensor $T_{ij}$ is computed using Eq. 17, which accounts for all interparticle contact forces within a cell. From this, we extract the normal and shear stress components relevant to the local material response.
The variables used in these equations are defined as follows: $x_{i}$ and $X_{j}$ are the deformed and reference coordinates, respectively; $x_{0i}$ is a constant translation vector; $v_{i}$ and $v^{0}_{i}$ are the components of the instantaneous and constant velocity vectors respectively; $L_{ij}$ and $L_{ji}$ represent the components of the velocity gradient tensor and its transpose, respectively; $D_{ij}$ is the component of the symmetric deformation rate tensor; $V$ is the RVE cell volume; $N$ is the number of particles in the cell; $N_{l}$ is the number of contacts for particle $l$; $\Delta x_{i}^{(k,l)}$ is the $i$-th component of the vector from the center of particle $l$ to its $k$-th contact; and $F_{j}^{(k,l)}$ is the $j$-th component of the force on particle $l$ from its $k$-th contact.
@@
Using the coarse-grained stress tensor, $T_{ij}$ (Eq. 17), we compute the shear stress magnitude, $\tau$ and the effective normal stress, $\sigma^{\prime}_{n}=\sigma_{n}-p$ (with $p$ the pore pressure), and obtain the apparent friction, $\mu_{\mathrm{}}=\tau/\sigma^{\prime}_{n}$ and normalized pore pressure, $p/\sigma^{\prime}_{n}$. To explore the nature of failure and flow, dimensionless numbers, specifically the viscous and inertial numbers, are used to distinguish between flow regimes by comparing competing physical forces. The viscous number, $I_{v}$, quantifies the relative importance of viscous forces to confining pressure, as defined in Eq. 18. It is commonly applied in systems where fluid resistance dominates particle dynamics, such as dense suspensions or saturated soils. Larger values of $I_{v}$ correspond to regimes where viscous effects are more significant. In contrast, the inertial number, $I_{n}$, characterizes the ratio of inertial to confining forces in dry or rapid granular flows, where particle collisions and momentum exchange dominate, as given by Eq. 18. Higher values of $I_{n}$ indicate stronger inertial effects. In these expressions, $\eta$ is the dynamic viscosity of the fluid, $\dot{\gamma}$ is the shear strain rate, $P$ is the confining pressure, $d_{p}$ is the particle diameter, and $\rho_{s}$ is the particle density.
$I_{v}=\frac{\eta\dot{\gamma}}{P}\quad\text{and}\quad I_{n}=\frac{\dot{\gamma}\,d_{p}}{\sqrt{P/\rho_{s}}}$ (18)
Figure 2: Comparison of failure and runout dynamics using velocity for subaqueous and subaerial models with dense ($\phi$ = 0.579) and loose ($\phi$ = 0.444) configurations at two time steps: 0.04 s and 0.4 s. Subaqueous model exhibit slower failure and reduced runout distances compared to subaerial model, due to particle–fluid interactions. Panels (a) and (b) show the dense subaqueous model with a minimal runout of 0.054 m (61% shorter than the maximum), while panels (c) and (d) represent the loose subaqueous model, which displays greater displacement and a more extended runout of 0.062 m (59% shorter). Panels (e) and (f) display the dense the subaerial model, characterized by abrupt failure with a runout of 0.15 m (maximum), and panels (g) and (h) illustrate the loose subaerial model, which also exhibits a runout of 0.15 m (maximum).
@@ III Mechanism of failure
The failure mechanisms of granular columns differ significantly between subaerial and subaqueous conditions, with the initial packing fraction playing a crucial role in these behaviors. Subaqueous systems exhibit reduced runout distances and slower failure for both dense and loose configurations due to particle–fluid interactions, with dense configuration being the most stable, a result consistent with previous studies Zhang2020; Montell2023. In contrast, subaerial systems fail more rapidly and exhibit greater sensitivity to packing fraction https://doi.org/10.48550/arxiv.1512.05180. Loose configurations undergo more pronounced and extended collapses, while dense configurations fail abruptly but with limited spreading. Fig. 2 compares velocity profiles at two time steps (0.04 s and 0.4 s) to illustrate the contrasting failure and runout behaviors across dense and loose configurations in both environments. In subaqueous systems, particle–fluid interactions, such as drag and pore pressure feedback, moderate failure dynamics, leading to slower progression and reduced energy dissipation compared to subaerial cases, although final runout distances may be similar in some cases Zhang2020.
@@
The dense subaqueous configuration (Fig. 2(a) and Fig. 2(b)) shows minimal runout, with a maximum of 0.054 m, which is approximately 61% shorter than the maximum runout observed in subaerial conditions. The loose subaqueous configuration (Fig. 2(c) and Fig. 2(d)) exhibits a slightly longer runout of 0.062 m, yet still 59% shorter than its subaerial counterpart. These reductions underscore the dampening influence of fluid on collapse extent. As shearing initiates, dilation occurs due to interlocking grains pushing against one another, increasing the pore pressure and mobilizing peak shear resistance. This interparticle interlocking elevates the internal friction angle and temporarily raises shear strength Montell2023. Following this peak, the friction angle decreases, and failure progresses gradually with limited velocity magnitudes throughout the collapse.
Figure 3: (a) Front-position evolution for a dense subaqueous column collapse. Orange line: CFD–DEM ($\phi=0.579$, $A=0.80$, 3D, laminar). Squares: Rondon et al. (2011) flume data ($\phi=0.60$, $A\approx 0.67$, quasi-2D). Both show rapid advance followed by a plateau; offsets reflect geometry and sidewall effects. (b) Normalized runout, $L/L_{i}$ for dense subaerial and subaqueous MFIX–DEM runs compared with Rondon data. Our subaqueous dense result ($L_{f}/L_{i}=2.92$ at $\phi=0.579$ and $3.21$ at $\phi=0.444$) falls within the experimental envelope ($L_{f}/L_{i}=2.05$–3.08 for $\phi=0.55$–0.60). Subaerial values ($L_{f}/L_{i}=3.69$ at $\phi=0.579$ and $4.01$ at $\phi=0.444$) are much larger than quasi-2D flumes ($L_{f}/L_{i}=2.6$ at $\phi=0.60$), reflecting reduced sidewall dissipation in 3D setups Jop2005. The loose case ($\phi=0.444$) lies below experimental packing ranges and densifies rapidly, preventing pore-pressure-driven mobility. Importantly, the relative ordering (subaerial >> subaqueous) remains consistent with experiments Rondon2011; Polania2024.
@@
Figure 4: Depiction of shear zone development and pore pressure evolution at two time steps (0.1 s and 0.8 s) for dense ($\phi$=0.579) and loose ($\phi$=0.444) subaqueous models. Panels (a) and (b) show shear zones, defined by zero volumetric strain within the critical state framework Wood1991, for the dense configuration at the two respective time steps, while panels (e) and (f) show the corresponding results for the loose configuration. Panels (c) and (d) illustrate pore pressure changes in the dense configuration, which begins with lower initial values that increase over time; in contrast, panels (g) and (h) show the loose configuration, which starts with higher initial pore pressures that decrease with continued deformation.